Simplify:
-6(bc^4)^-3
Please show me step by step to get the answer. thanks

Answer by stanbon(75887)

Simplify:
-6(bc^4)^-3
---
= -6(1/bc^4)^3
----
= -6(1/b^3c^12)
---
= -6/(b^3c^12)
====================
Cheers,
Stan H.